OpenStudy (anonymous):

(x^2-9)/(x^3+3x^2) divide by (x^2+x-12)/(x^3+9x^2+20x) help ASAP!

OpenStudy (anonymous):

when?

OpenStudy (anonymous):

\[ \frac{\left(x^2-9\right) \left(x^3+9 x^2+20 x\right)}{\left(x^2+x-12\right) \left(x^3+3 x^2\right)}=\frac{(x-3) (x+3) x (x+4) (x+5)}{(x-3) (x+4) x^2 (x+3)}=\frac{x+5}{x} \]

OpenStudy (anonymous):

merci

OpenStudy (anonymous):

Avec plaisir